1. Let U = {l, 2, 3, 4, 5, 6, 7, 8, 9, and 10} be a universal set. Let A, B, C such that A= {l, 3, 4, 8},
B = {2, 3, 4, 5, 9, 10}, and C = {3, 5, 7, 9, 10}. Use bit representations
for A, B, and C together with UNION,
INTER, DIFF, and COMP to find the bit representation for the following:
(a) AU B
(b) An B n C
(f)) (AU C) n B
(d) (A - B) UC
(e) An (B - (C n B))
(f) A - (B - C)
(g) (AU B) U (C - B)
1
Expert's answer
2020-09-10T18:41:37-0400

"A=\{1,3,4,8\}"

"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"


"B=\{2,3,4,5,9,10\}"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"


"C=\{3,5,7,9,10\}"

"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

Table for the Bit Operators OR,AND, and XOR


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c :c}
 x & y & x\lor y & x\land y & x\oplus y \\ \hline
 0 & 0 & 0 & 0 & 0 \\
 \hdashline 0 & 1 & 1 & 0 & 1 
 \\
 \hdashline 1 & 0 & 1 & 0 & 1 
 \\
 \hdashline 1 & 1 & 1 & 1 & 0 
 \\\end{array}"


a) "A\cup B"

"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"1\ 1\ 1\ 1\ 1\ 0\ 0\ 1\ 1\ 1" bitwise "OR"

"A\cup B=\{1,2,3,4,5,8,9,10\}"


b) "A\cap B\cap C"

"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 0\ 1\ 1\ 0\ 0\ 0\ 0\ 0\ 0" bitwise "AND"


"0\ 0\ 1\ 1\ 0\ 0\ 0\ 0\ 0\ 0"

"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

"0\ 0\ 1\ 0\ 0\ 0\ 0\ 0\ 0\ 0" bitwise "AND"

"A\cap B\cap C=\{3\}"


c) "(A\cup C)\cap B"

"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

"1\ 0\ 1\ 1\ 1\ 0\ 1\ 1\ 1\ 1" bitwise "OR"


"1\ 0\ 1\ 1\ 1\ 0\ 1\ 1\ 1\ 1"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 0\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1" bitwise "AND"

"(A\cup C)\cap B=\{3,4,5,9,10\}"


d) "(A-B)\cup C"

"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"1\ 0\ 0\ 0\ 0\ 0\ 0\ 1\ 0\ 0" bitwise "A-B"


"1\ 0\ 0\ 0\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

"1\ 0\ 1\ 0\ 1\ 0\ 1\ 1\ 1\ 1" bitwise "OR"

"(A-B)\cup C=\{1,3,5,7,8,9,10\}"


e) "A\cap (B-(C\cap B))"

"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 0\ 1\ 0\ 1\ 0\ 0\ 0\ 1\ 1" bitwise "AND"


"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 0\ 1\ 0\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 1\ 0\ 1\ 0\ 0\ 0\ 0\ 0\ 0" bitwise "B-(C\cap B)"


"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 1\ 0\ 1\ 0\ 0\ 0\ 0\ 0\ 0"

"0\ 0\ 0\ 1\ 0\ 0\ 0\ 0\ 0\ 0" bitwise "AND"

"A\cap (B-(C\cap B))=\{4\}"


f) "A-(B-C)"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

"0\ 1\ 0\ 1\ 0\ 0\ 0\ 0\ 0\ 0" bitwise "B-C"


"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 1\ 0\ 1\ 0\ 0\ 0\ 0\ 0\ 0"

"1\ 0\ 1\ 0\ 0\ 0\ 0\ 1\ 0\ 0" bitwise "A-(B-C)"

"A-(B-C)=\{1,3,8\}"


g) "(A\cup B)\cup(C-B)"

"1\ 0\ 1\ 1\ 0\ 0\ 0\ 1\ 0\ 0"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"1\ 1\ 1\ 1\ 1\ 0\ 0\ 1\ 1\ 1" bitwise "OR"


"0\ 0\ 1\ 0\ 1\ 0\ 1\ 0\ 1\ 1"

"0\ 1\ 1\ 1\ 1\ 0\ 0\ 0\ 1\ 1"

"0\ 0\ 0\ 0\ 0\ 0\ 1\ 0\ 0\ 0" bitwise "C-B"


"1\ 1\ 1\ 1\ 1\ 0\ 0\ 1\ 1\ 1"

"0\ 0\ 0\ 0\ 0\ 0\ 1\ 0\ 0\ 0"

"1\ 1\ 1\ 1\ 1\ 0\ 1\ 1\ 1\ 1" bitwise "OR"

"(A\cup B)\cup(C-B)=\{1,2,3,4,5,7,8,9,10\}"